# AppReferences

### v1

@@ Abstract
Principal stratification provides a causal inference framework that allows adjustment for confounded post-treatment variables when comparing treatments. Although the literature has focused mainly on binary post-treatment variables, there is a growing interest in principal stratification involving continuous post-treatment variables. However, characterizing the latent principal strata with a continuous post-treatment presents a significant challenge, which is further complicated in observational studies where the treatment is not randomized. In this paper, we introduce the Confounders-Aware SHared atoms BAyesian mixture (CASBAH), a novel approach for principal stratification with continuous post-treatment variables that can be directly applied to observational studies. CASBAH leverages a dependent Dirichlet process—utilizing shared atoms across treatment levels—to effectively control for measured confounders and facilitate information sharing between treatment groups in the identification of principal strata membership. CASBAH also offers a comprehensive quantification of uncertainty surrounding the membership of the principal strata. Through Monte Carlo simulations, we show that the proposed methodology has excellent performance in characterizing the latent principal strata and estimating the effects of treatment on post-treatment variables and outcomes. Finally, CASBAH is applied to a case study in which we estimate the causal effects of US national air quality regulations on pollution levels and health outcomes.
Keywords: Causal Inference; Dependent Dirichlet Process; Environmental Health; Principal Strata; Shared Atoms Mixture Model
@@ 1 Introduction
Principal stratification provides an effective causal inference approach to deal with post-treatment variables (Frangakis and Rubin, 2002). A post-treatment variable is defined as a variable that is potentially affected by treatment and also affects the primary potential outcomes. Thus, a post-treatment variable can be seen as an intermediate variable between the treatment and the outcome of interest. Post-treatment variables may pose inferential challenges for the estimation of treatment effects but also provide a useful approach to conceptualizing the mechanisms/pathways through which the treatment affects the outcome through the intermediate variable.
@@
Principal stratification with respect to the post-treatment variable—also known as intermediate variable—involves categorizing subjects into different groups, known as principal strata, based on the joint potential values of that post-treatment variable at various treatment levels (Antonelli et al., 2023). Principal strata consist of units that share similar potential values of the post-treatment—which is unaffected by treatment—thus representing an inherent latent characteristic of the units.
@@
Since the principal strata are, by definition, not influenced by the assignment of treatment, comparisons of potential outcomes at different treatment levels within a principal stratum represent well-defined causal effects (Rubin, 1974). These specific comparisons are referred to as principal causal effects. Researchers have often referred to the dissociative effect as the effect of treatment on the outcome for the stratum that includes units whose potential post-treatment variables do not notably change between different treatment levels, while labeling as the associative effect the effect of the treatment on the outcome for the stratum that includes units whose post-treatment variable do notably change between the different treatment levels (Frangakis and Rubin, 2002; Mealli and Mattei, 2012).
@@
Most of the literature on principal stratification and principal causal effects has focused on the case of binary treatment and the binary post-treatment variable (see, e.g., Angrist et al., 1996). This implies that researchers must deal with at most four combinations of these two variables (see the broad literature for binary or discrete post-treatment variables—e.g., Cheng and Small, 2006; Imai, 2008; Ding et al., 2011; Mattei and Mealli, 2011; Mealli and Pacini, 2013; Mealli et al., 2016; Jiang et al., 2016, 2022; Mattei et al., 2024). Considering continuous post-treatment variables leads to a potentially infinite number of potential principal strata, which, in turn, makes inference and interpretation very challenging. Various solutions have been proposed in the causal inference literature to deal with continuous post-treatments. Generally, it has been proposed to: (i) dichotomize the post-treatment variable (Sjölander et al., 2009; Jiang et al., 2022), or model the continuous post-treatment variable conditional to binary treatment using (ii) fully parametric or semiparametric modeling (Jin and Rubin, 2008; Conlon et al., 2014; Lu et al., 2023), or (iii) nonparametric modeling (Schwartz et al., 2011). Recently, Antonelli et al. (2023) has proposed an extension of nonparametric modeling in the case of continuous treatment using a Bayesian methodology.
@@
Although these methods provide potential solutions to the challenges of continuous post-treatment variables, they often require strong assumptions or subjective choices. On the one hand, dichotomization of the post-treatment variable requires the choice of the threshold on the post-treatment variable (e.g., Zigler et al., 2018; Antonelli et al., 2023; Lu et al., 2023). Although this choice might be informed by the relevant subject matter literature, it entails a level of arbitrability and requires extensive checking the robustness of the results with respect to the chosen thresholds.
@@
The method proposed in this paper extends the exciting work of Schwartz et al. (2011). The authors model a continuous post-treatment outcomes using a Dirichlet process mixture (DPM) method (Ferguson, 1973). We build upon it to allow for a novel and more interpretable characterization of the principal strata that overcomes some challenges. Firstly, our goal is to foster better information sharing between the marginal distribution of potential post-treatment variables, so to enhance the ability to uncover the principal strata and provide a more interpretable definition of them. Secondly, while the DPM proposed by Schwartz et al. (2011) adheres to the standard stick-breaking definition by Sethuraman (1994), our approach extends this by incorporating covariates into the weights of the mixture model. This adjustment improves the imputation of missing potential outcomes for the post-treatment, consequently enhancing the estimation performance of treatment effects.
@@
To do so, we introduce a novel Bayesian approach to principal stratification. Specifically, we introduce a Confounders-Aware SHared-atoms BAyesian mixture model (denoted by the acronym CASBAH, with the intentional alteration of the letter “H") endowed with innovative estimands for principal causal effects. CASBAH exploits the dependent Dirichlet process (Quintana et al., 2022) to establish a highly flexible distribution for the potential post-treatment variables given the confounders and the treatment variable, similar to the recent proposal of Zorzetto et al. (2024) in the context of heterogeneity in the treatment effects.
@@
CASBAH improves the current literature on nonparametric modeling for principal stratification in many respects. Firstly, it allows for the inclusion of the confounders in the weights of the infinite Bayesian mixture model, which provides a way to control for measured confounding and a more precise estimation of the causal effects. Secondly, the proposed model—which leverages shared random kernel’s atoms—allows us to identify associative/dissociative strata borrowing information across the treatment levels. Hence, we do not have to choose subjective criteria to define the principal strata, but CASBAH can, simultaneously, (i) identify the principal strata, according to the nature of the observed data, and (ii) impute the missing post-treatment variables and outcomes. Quantification of the uncertainty of the membership of the principal strata and estimation of the principal causal effects are inherent in CASBAH.
@@
This paper is organized as follows. In Section 2, we introduce the framework and the definition of the novel principal causal estimands. In Section 3, we propose CASBAH and how its characteristics allow us to identify the principal strata using a data-driven method. Theoretical prior properties provides insights on prior elicitation. An efficient Gibbs sampling algorithm to sample from the posterior distributions and estimate the model is also derived. Section 4 illustrates, via simulations, the ability of the proposed model to impute missing post-treatment outcomes, discover the main strata, and estimate the principal causal effects. In Section 5, we estimate the principal causal effects of air pollution policies on the mortality rate in the eastern United States (Zigler et al., 2018), where long-term fine particles pm2.5 are the post-treatment variable. In Section 6, we conclude the paper with a discussion and avenues for future work.
@@ 2 Setup
We assume that we observe $n$ independent units. For each unit $i\in\{1,\dots,n\}$, let $\mathbf{X}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{P}$ the set of observed covariates, $T_{i}\in\{0,1\}$ the observed (binary) treatment, $P_{i}\in\mathcal{P}\subseteq\mathbb{R}$ the post-treatment variable, and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ the response. According to Rubin’s Causal Model (Rubin, 1974), for each unit $i$, the potential outcome of the post-treatment variable is defined as $\{P_{i}(0),P_{i}(1)\}\in{\mathds{R}}^{2}$ and the potential outcome of the response is defined as $\{Y_{i}(0),Y_{i}(1)\}\in{\mathds{R}}^{2}$, for $i=1,\dots,n$. The vector $\{P_{i}(0),P_{i}(1)\}$ represents the collection of the two potential outcomes of the post-treatment variable, when the unit $i$ is assigned to the control or the treatment group, respectively. Similarly, the vector $\{Y_{i}(0),Y_{i}(1)\}$ represents the collection of the two potential outcomes for the response variable, with $Y_{i}(0)$ when the unit $i$ is assigned to the control group and $Y_{i}(1)$ when assigned to the treatment group.
@@ 2.1 Identifying Assumptions
Following the principal stratification framework (Frangakis and Rubin, 2002; VanderWeele, 2011; Mealli and Mattei, 2012; Feller et al., 2017; Ding et al., 2017; Lu et al., 2023), the principal strata are not affected by treatment assignment; therefore, a principal stratification can be used as any unit classification to define meaningful causal estimates conditional on the principal strata and discover the heterogeneous treatment effect (Mealli and Mattei, 2012).
As highlighted in the Introduction, the definition of the dissociative and associative strata is of particular interest. Formally, a dissociative stratum includes units where the treatment does not change the post-treatment variable, that is, $P_{i}(0)$ is the same as $P_{i}(1)$. In contrast, the associative stratum includes units for which treatment substantially affects the post-treatment variable, increasing it—positive associative stratum—or decreasing it—negative associative stratum.
@@
Assumption 1 (Stable Unit Treatment Value Assumption)
$Y_{i}(T_{i})=Y_{i}(T_{1},T_{2},\cdots,T_{i},\cdots,T_{n}),\quad Y_{i}(T_{i})=Y_{i},\mbox{ for }i=1,\dots,n;$
$P_{i}(T_{i})=P_{i}(T_{1},T_{2},\cdots,T_{i},\cdots,T_{n}),\quad P_{i}(T_{i})=P_{i},\mbox{ for }i=1,\dots,n.$
The Stable Unit Treatment Value Assumption (SUTVA) enforces that each unit’s outcome is a function of its treatment only. This is a combination of (i) no interference between units, that is, the potential outcome and the potential values of the post-treatment variable of the unit $i$ do not depend on the treatment applied to other units, and (ii) consistency, that is, no different versions of the treatment levels assigned to each unit (Rubin, 1986). Under the principal stratification framework, this assumption is invoked for both the outcome variable and the post-treatment variable.
In practice, under SUTVA, for $i=1,\dots,n$, we observe only $p_{i}\in{\mathds{R}}$ and $y_{i}\in{\mathds{R}}$, that is, the realization of the random variable $P_{i}$ and $Y_{i}$, respectively, and defined as $P_{i}=(1-T_{i})\cdot P_{i}(0)+T_{i}\cdot P_{i}(1)$ and $Y_{i}=(1-T_{i})\cdot Y_{i}(0)+T_{i}\cdot Y_{i}(1)$.
Assumption 2 (Strongly Ignorable Treatment Assignment)
Given the observed covariates ${x}_{i}$,
$\{\bigl{(}Y_{i}(1),Y_{i}(0)\bigl{)},\bigl{(}P_{i}(0),P_{i}(1)\bigl{)}\}\perp\!\!\!\perp T_{i}\mid X_{i},$
@@
where $\mathcal{X}$ is the features’ space.
This assumption states that: (i) for each unit, the potential outcome and the potential values of the post-treatment variable are independent of the treatment conditional on the set of covariates $X_{i}$; (ii) all units have a positive chance of receiving the treatment.
@@ 3 Confounders-Aware Shared-atoms Bayesian Mixture Model
Following the Bayesian paradigm and invoking the De Finetti’s theorem (Rubin, 1978), the joint probability distribution of the involved variables is unit exchangeable and it is defined as
@@
Following the assumptions in the previous section, the conditional probability of the treatment variable can be written as $\Pr(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)=\Pr(T_{i}\mid X_{i},\theta)$. Moreover, we condition the empirical distribution of covariates so that $\Pr(X_{i}\mid\theta)=\Pr(X_{i})$.
However, the remaining two probabilities in equation (1), have to be modeled: (i) the distribution of potential outcomes of the response conditional on the potential outcome of the post-treatment variable and covariates, and (ii) the distribution of the potential outcome of the post-treatment variable conditional on the covariates.
We focus our attention on the latter, that is, the distribution of the potential outcome of the post-treatment variable conditional on the covariates for which we propose a novel Bayesian nonparametric approach. As our primary focus lies in the post-treatment variable distribution, for the sake of clarity in our discussion, we make the simplifying assumption of employing a more straightforward parametric model for the conditional response distribution following the settings of Schwartz et al. (2011). Although the outcome model is assumed to be a linear regression model, it can also be generalized to a more complex and flexible model.
For the post-treatment variable, we previously exploited a dependent nonparametric mixture, following the approach adopted in the previous section of this paper and more broadly the dependent Dirichlet process (MacEachern, 2000; Barrientos et al., 2012; Quintana et al., 2022). In particular, our model also shares some similarities with the hierarchical Dirichlet process (Teh et al., 2004, 2006; Teh and Jordan, 2009) and the more recent common atom model of Denti et al. (2023).
Specifically, we assume for each $i=1,\dots,n$:
$\{P_{i}\mid{x}_{i},t\}$ $\sim f^{(t)}(\cdot\mid{x}_{i}),\mbox{ for }t=\{0,1\},$
@@
$\{G_{{x}_{i}}^{(0)},G_{{x}_{i}}^{(1)}\}$ $\sim\Pi,$ (2)
where ${\mathcal{K}}(\cdot;x,\psi)$ is a continuous density function, for every $\psi\in\Psi$, and $G_{{x}_{i}}^{(t)}$ is a random probability measure depending on the confounders $x_{i}$ associated to an observation assigned to treatment level $t$. In particular, we have two treatment levels $t=\{0,1\}$, therefore, the distributions belonging to $G_{{x}_{i}}^{(0)}$ and $G_{{x}_{i}}^{(1)}$ are characterized by specific weights assigned to a common set of atoms induced by the common process $\Pi$.
Following the characterization of the random probability measures we can write:
$G_{{x}_{i}}^{(t)}=\sum_{l\geq 1}\pi_{l}^{(t)}({x}_{i})\delta_{\mathbb{\psi}_{l}},$ (3)
where the sequences $\{\pi_{l}^{(t)}(x_{i})\}_{l\geq 1}$ for $t=\{0,1\}$ represent infinite sequences of random weights, and $\{\psi_{l}\}_{l\geq 1}$ is an infinite sequence of random kernel’s parameters—independent and identically distributed from a base measure $H$—shared among potential outcome of post-treatment variable under both the treatment level $t$.
Furthermore, the sequences of weights are defined through a stick-breaking representation (Sethuraman, 1994),
$\pi_{l}^{(t)}({x}_{i})$ $=\omega_{l}^{(t)}({x}_{i})\prod_{r<l}\{1-\omega_{r}^{(t)}({x}_{i})\};$ (4)
where $\{\omega_{l}^{(t)}({x})\}_{l\geq 1}$ , for $t=\{0,1\}$ are $[0,1]$-valued independent stochastic processes.
The discrete nature of the random probability measure $G_{{x}_{i}}^{(t)}$, for $t=\{0,1\}$, allows us to introduce the latent categorical variables $S_{i}^{(t)}$, for $t=\{0,1\}$, describing the groups of units defined by heterogeneity in $P_{i}(t)$. Assuming $\mathds{P}\{S_{i}^{(t)}=l\}=\pi_{l}^{(t)}({x}_{i})$, we can write model in (2), exploiting conditioning on $S_{i}^{(t)}$, as
$\{P_{i}|{x}_{i},t,\psi,S_{i}^{(t)}=l\}\sim{\mathcal{K}}(\cdot\mid{x}_{i},\psi_{l}),\quad\psi_{l}\sim H.$
@@
Among the plethora of dependent non-parametric processes (see the recent review by Quintana et al., 2022, for a detailed description), we focused on the dependent probit stick breaking process for its success in applications, its good theoretical properties and ease of computation (Rodriguez and Dunson, 2011). Consistently with this, we specify:
@@
As commonly done, we assume the kernel $\mathcal{K}$ to be a Gaussian, so that model (2)–(3) specifies to
$\{P_{i}(t)|{x}_{i},S_{i}^{(t)}=l,\eta,\sigma\}\sim\mathcal{N}(\eta_{l},\sigma_{l}^{2}).$ (6)
@@
where InvGamma($\gamma_{1},\gamma_{2}$) represents the inverse gamma distribution with the shape parameter $\gamma_{1}\in\mathbb{R}^{+}$ and the scale parameter $\gamma_{2}\in\mathbb{R}^{+}$, and mean equal to $\frac{\gamma_{2}}{\gamma_{1}-1}$ and variance $\frac{\gamma_{2}^{2}}{(\gamma_{1}-1)^{2}(\gamma_{1}-2)}$.

### v2

@@ Abstract
Principal stratification provides a causal inference framework for investigating treatment effects in the presence of a post-treatment variable. Principal strata play a key role in characterizing the treatment effect by identifying groups of units with the same or similar values for the potential post-treatment variable under both treatment levels. The literature has focused mainly on binary post-treatment variables. Few papers considered continuous post-treatment variables. In the presence of a continuous post-treatment, a challenge is how to identify and characterize meaningful coarsening of the latent principal strata that lead to interpretable principal causal effects. This paper introduces the confounders-aware shared-atom Bayesian mixture, a novel approach for principal stratification with binary treatment and continuous post-treatment variables. Our method leverages Bayesian nonparametric priors with an innovative hierarchical structure for the potential post-treatment variable that overcomes some of the limitations of previous works. Specifically, the novel features of our method allow for (i) identifying coarsened principal strata through a data-adaptive approach and (ii) providing a comprehensive quantification of the uncertainty surrounding stratum membership. Through Monte Carlo simulations, we show that the proposed methodology performs better than existing methods in characterizing the principal strata and estimating principal effects of the treatment. Finally, our proposed model is applied to a case study in which we estimate the causal effects of US national air quality regulations on pollution levels and health outcomes.
keywords:
Bayesian Causal Inference; Principal Stratification; Shared Atom Mixture Model.
@@ 1 Introduction
Principal stratification is a well-established causal inference approach that accounts for post-treatment variables (Frangakis & Rubin, 2002). A post-treatment variable is defined as a variable that is potentially affected by treatment and it is also associated with the primary potential outcomes. Thus, a post-treatment variable can be seen as an intermediate variable between the treatment and the outcome of interest. Post-treatment variables may pose inferential challenges for the estimation of treatment effects, but also provide a useful approach to conceptualizing the mechanisms/pathways through which the treatment affects the outcome through the intermediate variable.
@@
Principal stratification with respect to the post-treatment variable, involves stratifying units into different groups, known as principal strata. Principal strata consist of units that share the same or similar potential values of the post-treatment variable under all possible treatment levels: by definition, principal strata are not affected by the treatment and thus represent an inherent latent characteristic of the units.
@@
Since the assignment of treatment does not influence the principal strata, comparisons of potential outcomes at different treatment levels within a principal stratum represent well-defined causal effects (Rubin, 1974). These specific comparisons are referred to as principal causal effects. Researchers have often referred to the dissociative (and associative) effect as the effect of treatment on the outcome for the stratum that includes units whose potential post-treatment outcomes do not change (or change) under different treatment levels (Frangakis & Rubin, 2002; Mealli & Mattei, 2012).
@@
Most of the literature on principal stratification and principal causal effects has focused on the case of binary treatment and the binary post-treatment variable (see, e.g., Angrist et al., 1996). This implies that researchers deal with at most four principal strata that are the possible combinations of values of these two variables (see the broad literature for binary or discrete post-treatment variables, e.g., Cheng & Small, 2006; Imai, 2008; Ding et al., 2011; Mattei & Mealli, 2011; Mealli & Pacini, 2013; Mealli et al., 2016; Jiang et al., 2016, 2022; Mattei et al., 2024). Considering continuous post-treatment variables leads to a potentially infinite number of principal strata, which, in turn, makes inference and interpretation very challenging. Various solutions have been proposed in the causal inference literature. Proposals include: (i) dichotomize the post-treatment variable (Sjölander et al., 2009; Jiang et al., 2022), or model the continuous post-treatment variable conditional on the binary treatment using (ii) fully parametric or semiparametric modeling (Jin & Rubin, 2008; Conlon et al., 2014; Lu et al., 2023), or (iii) nonparametric modeling (Schwartz et al., 2011).
@@
Although these methods provide potential solutions to the challenges of continuous post-treatment variables, they often require strong assumptions or subjective choices. On the one hand, dichotomization of the post-treatment variable requires the choice of the threshold on the post-treatment variable (e.g., Zigler et al., 2012; Antonelli et al., 2023). Although this choice might be informed by the relevant subject matter literature, it entails a level of arbitrariness, and requires extensive checking of the robustness of the results with respect to the chosen thresholds.
@@
Our proposed method introduces a novel and more interpretable approach to characterizing the principal strata, addressing some critical challenges in this framework. While, similarly Schwartz et al. (2011) modeled the continuous post-treatment variable using a Dirichlet process mixture (Ferguson, 1973), our approach advances this framework in two significant ways. First, we foster information sharing between the marginal distribution of potential post-treatment variables to enhance the ability to uncover the principal strata with a data-adaptive approach and provide a more interpretable definition of them. Second, while the Dirichlet process mixture proposed by Schwartz et al. (2011) adheres to the standard stick-breaking definition by Sethuraman (1994), our approach extends this by incorporating covariates into the weights of the mixture model. This adjustment significantly improves the imputation of missing potential post-treatment outcomes and enables strata identification without requiring predefined assumptions, consequently enhancing the estimation performance of treatment effects.
@@
To do so, we introduce a novel Bayesian approach to principal stratification. Specifically, we introduce a confounders-aware shared-atoms Bayesian mixture model endowed with innovative estimands for principal causal effects. Our model exploits the dependent Dirichlet process (Quintana et al., 2022) to establish a highly flexible distribution for the potential post-treatment outcomes given the confounders and the treatment variable, similar to the recent proposal of Zorzetto et al. (2024) in the context of heterogeneity in the treatment effects.
@@
Our proposed approach improves the current literature on nonparametric modeling for principal stratification in many respects. Firstly, it allows for the inclusion of the confounders in the weights of the infinite Bayesian mixture model, which provides a way to control for measured confounding and a more precise estimation of the causal effects. Secondly, the proposed model, which leverages shared random kernels’ atoms, allows us to identify associative/dissociative strata borrowing information across the treatment levels. Hence, our approach eliminates the need for subjective criteria to cluster principal strata. Instead, our model simultaneously (i) identifies the principal strata, according to the nature of the observed data, and (ii) imputes the missing post-treatment variables and outcomes. Moreover, our framework naturally incorporates uncertainty quantification for principal strata membership while providing robust estimation of principal causal effects.
@@
This paper is organized as follows. In Section 2, we introduce the framework and the definition of the novel principal causal estimands. In Section 3, we propose our model and detail how it allows us to identify meaningful coarsening of the principal strata using a data-adaptive method. Theoretical prior properties provide insight into prior elicitation. An efficient Gibbs sampling algorithm to sample from the posterior distributions and estimate the model is also derived. Section 4 illustrates, via simulations, the ability of the proposed model to impute missing post-treatment outcomes, discover the main strata, and estimate the principal causal effects. In Section 5, we estimate the principal causal effects of air pollution policies on the mortality rate in the eastern United States (Zigler et al., 2018), where long-term fine particle exposure is the post-treatment variable. In Section 6, we conclude the paper with a discussion and avenues for future work.
@@ 2 Setup
We assume that we observe $n$ independent units. For each unit $i\in\{1,\dots,n\}$, let $\mathbf{X}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{P}$ the set of observed covariates, $T_{i}\in\{0,1\}$ the observed (binary) treatment, $P_{i}\in\mathcal{P}\subseteq\mathbb{R}$ the post-treatment variable, and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ the primary response. For each unit $i$, the potential post-treatment outcomes are defined as $\{P_{i}(0),P_{i}(1)\}\in{\mathds{R}}^{2}$ and the potential primary outcomes are defined as $\{Y_{i}(0),Y_{i}(1)\}\in{\mathds{R}}^{2}$, for $i=1,\dots,n$. The vector $\{P_{i}(0),P_{i}(1)\}$ represents the collection of the two potential values of the post-treatment variable, when the unit $i$ is assigned to the control or the treatment group, respectively. Similarly, the vector $\{Y_{i}(0),Y_{i}(1)\}$ represents the collection of the two potential values for the response variable, with $Y_{i}(0)$ when the unit $i$ is assigned to the control group and $Y_{i}(1)$ when assigned to the treatment group.
@@ 2.1 Identifying Assumptions
Following the principal stratification framework (Frangakis & Rubin, 2002; VanderWeele, 2011; Mealli & Mattei, 2012; Feller et al., 2017; Ding et al., 2017; Lu et al., 2023), the principal strata are not affected by treatment assignment; therefore, a principal stratification can be used as any unit classification to define meaningful causal estimates conditional on the principal strata and discover the heterogeneous treatment effect (Mealli & Mattei, 2012).
As highlighted in the Introduction, the definition of the dissociative and associative strata is of particular interest. Formally, a dissociative stratum includes units where the treatment does not affect the post-treatment variable, that is, $P_{i}(0)$ is the same as $P_{i}(1)$. In contrast, the associative stratum includes units for which the treatment affects the post-treatment variable, yielding a positive effect, positive associative stratum, or a negative effect, negative associative stratum.
@@
{assumption}
[Stable Unit Treatment Value Assumption]
$Y_{i}(T_{1},T_{2},\cdots,T_{i},\cdots,T_{n})=Y_{i}(T_{i}),\quad Y_{i}(T_{i})=Y_{i},\mbox{ for }i=1,\dots,n;$
$P_{i}(T_{1},T_{2},\cdots,T_{i},\cdots,T_{n})=P_{i}(T_{i}),\quad P_{i}(T_{i})=P_{i},\mbox{ for }i=1,\dots,n.$
The Stable Unit Treatment Value Assumption (SUTVA) is a combination of (i) no interference between units, that is, the potential outcome and the potential values of the post-treatment variable of the unit $i$ do not depend on the treatment applied to other units, and (ii) consistency, that is, no different versions of the treatment levels assigned to each unit (Rubin, 1986). Under the principal stratification framework, this assumption is invoked for both the primary outcome variable and the post-treatment variable.
In practice, for $i=1,\dots,n$, we observe $p_{i}\in{\mathds{R}}$ and $y_{i}\in{\mathds{R}}$, that is, the realization of the random variables $P_{i}$ and $Y_{i}$, respectively, and defined as $P_{i}=(1-T_{i})\cdot P_{i}(0)+T_{i}\cdot P_{i}(1)$ and $Y_{i}=(1-T_{i})\cdot Y_{i}(0)+T_{i}\cdot Y_{i}(1)$.
{assumption}
[Strongly Ignorable Treatment Assignment] Given the observed covariates $\mathbf{x}_{i}$,
$\{Y_{i}(1),Y_{i}(0),P_{i}(0),P_{i}(1)\}\perp\!\!\!\perp T_{i}\mid X_{i},$
@@
where $\mathcal{X}$ is the features’ space. This assumption states that: (i) for each unit, the potential primary outcomes and the potential post-treatment variables are independent of the treatment conditional on the set of covariates $X_{i}$; (ii) all units have a positive chance of receiving the treatment.
@@ 3 Confounders-Aware Shared-atoms Bayesian Mixture Model
Following the Bayesian paradigm and invoking De Finetti’s theorem (Rubin, 1978), the joint probability distribution of the involved variables is unit exchangeable, and it is defined as
@@
Invoking Assumption 2, the conditional probability of the treatment variable can be written as $\Pr(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)=\Pr(T_{i}\mid X_{i},\theta)$. Moreover, we replace the conditional distribution of the covariates with the empirical distribution so that $\Pr(X_{i}\mid\theta)=\Pr(X_{i})$.
The remaining two probabilities in equation (1) must be modeled: (i) the distribution of potential primary outcomes conditional on the potential post-treatment variables and covariates, and (ii) the distribution of the potential post-treatment outcomes conditional on the covariates.
We focus our attention on the latter, that is, the distribution of the potential post-treatment outcomes conditional on the covariates for which we propose a novel Bayesian nonparametric approach. As our primary focus lies in modeling the post-treatment variable distribution, for the sake of clarity in our discussion, we employ a parametric model for the conditional response distribution following the settings of Schwartz et al. (2011). Note that, although the primary outcome model is assumed to be a linear regression model, it can also be generalized to a more complex and flexible model.
For modeling the post-treatment variable, we exploit a dependent nonparametric mixture, following the dependent Dirichlet process approach (MacEachern, 2000; Barrientos et al., 2012; Quintana et al., 2022; Denti et al., 2023). [^1] In particular, our model also shares some similarities with the hierarchical Dirichlet process (Teh et al., 2004, 2006; Teh & Jordan, 2009) and the recent common atom model of Denti et al. (2023). Specifically, we assume for each $i=1,\dots,n$:
$\{P_{i}\mid\mathbf{x}_{i},t\}$ $\sim f^{(t)}(\cdot\mid\mathbf{x}_{i}),\mbox{ for }t=\{0,1\},$
@@
$\{G_{\mathbf{x}_{i}}^{(0)},G_{\mathbf{x}_{i}}^{(1)}\}$ $\sim\Pi,$ (2)
where ${\mathcal{K}}(\cdot;x,\psi)$ is a continuous density function, for every $\psi\in\Psi$, and $G_{\mathbf{x}_{i}}^{(t)}$ is a random probability measure depending on the confounders $x_{i}$ associated to an observation assigned to treatment level $t$. For each treatment level $t=\{0,1\}$, the distributions of $G_{\mathbf{x}_{i}}^{(0)}$ and $G_{\mathbf{x}_{i}}^{(1)}$ are characterized by specific weights assigned to a common set of atoms induced by the common process $\Pi$.
Following the characterization of the random probability measures, we write:
$G_{\mathbf{x}_{i}}^{(t)}=\sum_{l\geq 1}\pi_{l}^{(t)}(\mathbf{x}_{i})\delta_{\mathbb{\psi}_{l}},$ (3)
where the sequences $\{\pi_{l}^{(t)}(x_{i})\}_{l\geq 1}$ for $t=\{0,1\}$ represent infinite sequences of random weights, and $\{\psi_{l}\}_{l\geq 1}$ is an infinite sequence of random kernel’s parameters, independent and identically distributed from a base measure $H$, shared among potential post-treatment outcomes under both treatment levels $t$.
The sequences of weights are defined through a stick-breaking representation (Sethuraman, 1994),
$\pi_{l}^{(t)}(\mathbf{x}_{i})$ $=\omega_{l}^{(t)}(\mathbf{x}_{i})\prod_{r<l}\{1-\omega_{r}^{(t)}(\mathbf{x}_{i})\};$ (4)
where $\{\omega_{l}^{(t)}(\mathbf{x})\}_{l\geq 1}$ , for $t=\{0,1\}$ are $[0,1]$-valued independent stochastic processes.
The discrete nature of the random probability measure $G_{\mathbf{x}_{i}}^{(t)}$, for $t=\{0,1\}$, allows us to introduce the latent categorical variables $S_{i}^{(t)}$, that describe the cluster allocations for each unit $i\in\{1,\dots,n\}$, whose clusters are defined by heterogeneous values for $P_{i}(t)$. Assuming $\mathds{P}\{S_{i}^{(t)}=l\}=\pi_{l}^{(t)}(\mathbf{x}_{i})$, we can write model in (2), exploiting conditioning on $S_{i}^{(t)}$, as
$\{P_{i}|\mathbf{x}_{i},t,\psi,S_{i}^{(t)}=l\}\sim{\mathcal{K}}(\cdot\mid\mathbf{x}_{i},\psi_{l}),\quad\psi_{l}\sim H.$
@@
Among the plethora of dependent non-parametric processes (see the recent review by Quintana et al., 2022, for a detailed description), we focused on the dependent probit stick-breaking process (Rodriguez & Dunson, 2011). Consistently with this, we specify:
@@
We assume the kernel $\mathcal{K}$ to be Gaussian, so that model (2)–(3) becomes
$\{P_{i}(t)\mid\mathbf{x}_{i},S_{i}^{(t)}=l,\eta,\sigma\}\sim\mathcal{N}(\eta_{l},\sigma_{l}^{2}).$ (6)
@@
where InvGamma($\gamma_{1},\gamma_{2}$) represents the inverse gamma distribution with the shape parameter $\gamma_{1}\in\mathbb{R}^{+}$ and the scale parameter $\gamma_{2}\in\mathbb{R}^{+}$, and mean equal to $\gamma_{2}/(\gamma_{1}-1)$ and variance $\gamma_{2}^{2}/\{(\gamma_{1}-1)^{2}(\gamma_{1}-2)\}$.